OpenStudy (anonymous):

Find antiderivative g(t) = 3 + t + t^2 / [sqrt(t) ]

OpenStudy (anonymous):

question edited^

OpenStudy (jziggy):

Well, firstly you can break it up into 3 parts, do you know how to find the antiderivative of the first 2 terms?

hartnn (hartnn):

just to clarify, is it (3+t+t^2) in the numerator or ? 3+t is separate and only t^2 in the numerator?

OpenStudy (anonymous):

first one

OpenStudy (jziggy):

ooh... that changes things

OpenStudy (jziggy):

a little

hartnn (hartnn):

thought so :)

OpenStudy (jziggy):

you can rewrite it with each term seperately over sqrt(t) and then simplify to get, 3t^(-1/2) + t^(1/2) + t^(3/2)

hartnn (hartnn):

its like separating out the numerator and then simplifying (a+b+c)/d = a/d+b/d+c/d

OpenStudy (jziggy):

Then you can just use the power rule for each individual term, the antiderivative of x^n = [x^(n+1)]/(n+1)